missrosejensen:

Which of the following is a solution to the equation x - = -? 2 1 -

Shadow:

This question is super incomplete

Shadow:

Image maybe?

missrosejensen:

Shadow:

\[x - \frac{ 5 }{ 7 } = -\frac{ 3 }{ 7 } \] We have two fractions with the same denominator, 7, so lets multiple the entire equation by 7 \[7(x - \frac{ 5 }{ 7 }) = (-\frac{ 3 }{ 7 }) \times 7\] We get: \[7x - 5 = -3\] Add 5 to both sides \[7x = 2\] \[x = \frac{ 2 }{ 7 }\]

missrosejensen:

thank you are you smart or what